OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y less than or greater to -2x + 8

OpenStudy (anonymous):

(0, -4) (5, -1) (7, 0) (9, 1)

OpenStudy (anonymous):

do you mean less then or equal ?

OpenStudy (anonymous):

if so, lets sub some points in and see which one satisfies the inequality. (0,-4) x = 0 and y = -4 y <= -2x + 8 -4 <= -2(0) + 8 -4 <= 8 (correct) This is correct. I was gonna do one as an example, but the first one happens to be the answer. Let me check the second one... (5,-1) x = 5 and y = -1 y <= -2x + 8 -1 <= -2(5) + 8 -1 <= -10 + 8 -1 <= -2 (false) -1 is greater then 2. anyways, do you understand how to do it now ? Just sub in your points and see which one is correct.